OpenStudy (anonymous):

The legs of an isosceles right triangle are 11 cm long. Find the length of the hypotenuse. Round to the nearest tenth if necessary. A. 15.6 cm B. 22 cm C. 7.8 cm D. 19.1 cm

OpenStudy (hoblos):

h^2 = 2(11)^2 = 242 h=sqrt(242) = 15.6 cm